# A Survey of Zero-Knowledge Proof Based Verifiable Machine Learning

### v1

@@ III-B1 ZKP based Verifiable Training
In 2021, Zhao et al. proposed VeriML [29], whose core idea is to make the training process retrievable to achieve verifiability in machine learning training. VeriML pre-stores the inputs and outputs of several iterations during the training process and commits to them, allowing the prover to retrieve specified iterations upon the verifier’s request and generate proofs of their computational processes. VeriML supports six typical machine learning models, including linear regression, logistic regression, neural networks, SVM, K-Means, and decision trees. Also, VeriML’s computational and communication costs are justified through numerous experiments. Experimental results demonstrate that the communication overhead of VeriML is related to the number of iterations and the number of data samples, and the computational overhead of VeriML is dominated by the batch size of data samples and increases linearly with the batch size.
@@
Zhang et al. introduced Hydra [30], a verifiable training protocol for neural networks built upon the GKR protocol. This system is specifically tailored for neural networks and builds upon the SafetyNets method—the first verifiable inference approach leveraging ZKP, which will be discussed later. In particular, Hydra uses SafetyNets to represent a neural network as an arithmetic circuit. To achieve verifiable training, Hydra introduces a subcircuit protocol and a interactive quantization algorithm. The execution process of Hydra’s subcircuit protocol can be summarized as follows: First, the large and deep circuit representing a neural network is partitioned into multiple smaller sub-circuits by depth. The prover then applies the GKR protocol to generate proofs for each sub-circuit. Once a proof for a sub-circuit is generated, it is immediately transmitted to the verifier for verification. Finally, all proofs corresponding to the sub-circuits are aggregated into a single final proof for the entire circuit, which is sent to the verifier for verification. This pipelined approach enables the proof-and-verification process to begin as soon as sub-circuits are uploaded, eliminating the need to process the entire circuit at once. By overlapping proof generation and verification, Hydra significantly enhances verification efficiency. Hydra’s interactive quantization algorithm begins by rounding the weights of the bottom layers of the neural network, freezing these quantized layers, and retraining the network. This process is then repeated layer by layer, progressively moving upward, until all layers have been quantized. This interactive quantization algorithm is executed for the each part of the neural network corresponding to each sub-circuit in the pipelined proof-and-verification process.
Figure 6: A timeline of representative existing ZKML studies.
@@
Huang et al. proposed zkMLaaS [31], a framework that employs a two-round challenge-response protocol and random sampling of model weights to generate proofs. This approach reduces the time cost of proof generation while ensuring the verifiable integrity of the training process. Prior to executing the challenge-response protocol, an honest third party (or the ML client) generates a key pair (PK,VK), distributing PK to the ML service provider (the prover) and VK to the ML client (the verifier). In the first round of the challenge-response protocol, the ML service provider submits commitments for all intermediate weights updated during each training iteration and data sampling epoch to the ML client. In the second round, the ML client randomly selects a subset of the intermediate weights, and the ML service provider is required to generate corresponding proofs for them. The zkMLaaS framework leverages zk-SNARKs and commitment schemes to ensure security. The binding property of the commitment scheme guarantees that the ML service provider cannot produce two different sets of weights corresponding to the same commitment. Furthermore, since the public key is generated by the client or a trusted third party, the service provider is restricted to generating proofs only for the circuits specified by the client. This ensures that the intermediate weights are derived solely from the correct model and data.
@@
zkDL [32] is an efficient ZKML system designed for the verifiable training of deep neural networks. To address the non-arithmetic nature of ReLU activation functions in neural networks, zkDL introduces the zkReLU protocol, which facilitates proving computations involving ReLU. The zkReLU protocol leverages auxiliary inputs to transform the nonlinear computations of ReLU into an equivalent linear form, enabling efficient handling of forward and backward propagations through ReLU. To ensure the validity of these auxiliary inputs, zkReLU employs the Pedersen commitment scheme and incorporates an anchoring mechanism that links the arithmetic operations within each layer to the auxiliary inputs. This design not only preserves the tensor structure inherent in deep neural networks but also significantly reduces computational redundancy by reusing verified commitments. For circuit design in the training of deep neural networks, zkDL proposes FAC4DNN (Flat Arithmetic Circuit for Deep Neural Networks), which parallelizes the traditionally sequential execution of neural network layers and training iterations. This approach effectively flattens the circuit, reducing its depth by a factor of $O(N)$, where $N$ is the product of the neural network depth and the number of training iterations. zkDL implements a three-step parallel proof process for FAC4DNN: (1) parallel execution of the GKR protocol for each layer, (2) parallel proof of inter-layer arithmetic relationships, and (3) parallel verification of auxiliary inputs. This design ensures that the proof size grows only by $O(\log L)$, where $L$ is the product of the neural network depth. Compared to traditional sequential proof generation, zkReLU’s parallel proof generation offers significant time efficiency. It shows in [32] that the proof generation of zkDL for the entire forward and backward propagation of the neural network training process to be completed within tens of seconds for neural networks with 10 millions parameters.
Grag et al. [33] proposed the concept of Zero-Knowledge Proof of Training (zkPoT), aiming to strike a balance between proof size and computation time. zkPoT is a novel and efficient zero-knowledge proof generation framework that combines two types of zero-knowledge proof techniques: arithmetic and non-arithmetic operations. By carefully integrating these techniques, zkPoT seeks to overcome the limitations of existing methods, ensuring that the proof size grows linearly with the number of the data in the dataset but remains independent of the number of data features. Moreover, this framework does not require the entire computation process to be stored in main memory (RAM) during proof generation, as it can load data from auxiliary memory when needed. Consequently, this approach imposes no fundamental memory limitations on the model size or dataset during training, making it more compatible with machine learning training tasks.
TABLE III: Classification of Existing ZKML Approaches
Research Works Time Application Verifiable Training Verifiable Testing Verifiable Inferring
@@
Since the zkPoT framework developed by Garg et al. [33] only supports basic machine learning algorithms, such as logistic regression, and does not extend to deep neural networks, Kasra et al. [49] proposed an efficient and concise zkPoT technique designed specifically for training DNN models using multiple small-batch gradient descent algorithms. Their work focuses on two key components: an optimized proof system for gradient descent iterations and a framework for efficient recursive proof composition across multiple iterations. First, Kasra et al. introduced an optimized GKR-style proof system (based on the sumcheck protocol) tailored for the gradient descent algorithm, enabling the verification of the correctness of computations in each iteration. Next, they developed a method to combine multiple proofs into a single, smaller proof, which can verify the correctness of results across multiple iterations, thereby achieving conciseness. Additionally, the authors proposed a general framework that integrates multiple proofs with polynomial commitments to produce a compact proof and commitment, significantly enhancing the efficiency of both proof generation and verification.
III-B2 ZKP based Verifiable Testing
@@
In 2020, Zhang et al. proposed a ZKML protocol for verifiable inference and accuracy testing of decision trees, named zkDT [36]. zkDT enables the owner of a decision tree model to convince others of the model’s inference results on data samples or its accuracy on public datasets, all without revealing any information about the model itself. The protocol leverages the Aurora protocol [53] as the ZKP backend due to Aurora’s fast proving time, which is a desirable feature for large decision trees. zkDT significantly improves the efficiency of proving time by transforming the computation of decision tree inference into a circuit of size $O(d+h)$, where $d$ is the length of the inference path on the tree, and $h$ is the number of data features. Notably, many testing data samples share common nodes on the decision tree. To exploit this, zkDT optimizes the proof and verification process for decision tree accuracy testing by validating all the nodes of the inference paths across all testing data in a single step, rather than validating the inference path of each sample individually. The implementation of zkDT demonstrates its practicality. For a decision tree with 23 levels and 1029 nodes, and a test dataset consisting of 5000 data samples with 54 features each, zkDT takes 250 seconds to generate a proof of size 287 KB for accuracy testing, and 15.6 seconds for verification.
Campanelli et al. [44] proposed a novel lookup argument and demonstrated how it can significantly improve upon the zkDT framework introduced in [36]. Lookup arguments allow one to prove that the elements of a committed vector originate from a larger committed table. This enables innovative approaches to reduce the prover complexity of general-purpose zk-SNARKs, particularly for implementing “non-arithmetic operations” such as range checks, XOR, and AND more efficiently. The authors make several key advancements in lookup arguments: (1) They extend vector lookups to matrix lookups, allowing proof that a committed matrix is a submatrix of a committed table. (2) They introduce a concept that ensures the privacy of both the sub-vector/sub-matrix and the table. (3) They propose new zero-knowledge lookup arguments, namely cq+, zkcq+, and cq++, which are more efficient than the recent work by Eagen, Fiore, and Gabizon, referred to as cq [54]. Furthermore, they present an application of fully zero-knowledge matrix lookup arguments to zkDT. Using matrix lookup arguments, one can commit to a decision tree by encoding it as a matrix, and to prove correct evaluation, the prover commits to the single row corresponding to the correct leaf. They can then prove, using the matrix lookup argument, that the committed row is indeed a leaf of the committed decision tree. Once the relevant row is isolated, the prover can then demonstrate that the input vector satisfies all the constraints outlined by that row. For statements involving multiple input vectors, rather than committing to a single row, the prover can commit to a matrix whose rows correspond to the entries of the leaves reached by the evaluations. Thanks to the efficiency of the matrix lookup argument, the prover’s time complexity remains independent of the size of the decision tree.
pvCNN [41] is a framework designed for the verifiable testing of convolutional neural networks (CNNs), developed by integrating fully homomorphic encryption, zk-SNARKs, and collaborative inference. The implementation of pvCNN can be summarized in three key steps. (1) pvCNN constructs a Quadratic Matrix Programs (QMP) based arithmetic circuit, which significantly reduces the number of multiplication gates required for convolution operations. This reduction minimizes the circuit size and enhances proof efficiency. Within this circuit, each multiplication gate represents the two-dimensional convolution operation between multiple filters and input data in a batch-processing manner. This design enables pvCNN to achieve zk-SNARK proof aggregation, allowing for the batch verification of multiple proofs. (2) pvCNN aggregates multiple proofs corresponding to test data provided by different testers for the same CNN model into a single proof. The validity of the aggregated proof guarantees the correctness of the original individual proofs, thereby improving verification efficiency. (3) The CNN model is divided into two parts: a private part, retained locally by the model developer, and a public part, outsourced to an external server. The private part processes test data encrypted via homomorphic encryption, as provided by the tester, and transmits its output to the public part to complete subsequent computations for CNN testing. This partitioning mechanism ensures privacy protection while enabling efficient inference verification. Through these steps, the pvCNN framework achieves efficient verification of neural network inference results from multiple testers while preserving data privacy. Experimental results demonstrate that, for high-dimensional matrix multiplication, pvCNN achieves a proof generation time that is approximately 13.9 times faster and a setup time that is 17.6 times faster than existing zk-SNARKs based on Quadratic Arithmetic Programs (QAP). Moreover, pvCNN overcomes the limitations of QAP-based zk-SNARKs in handling a bounded number of multiplication operations.
Feng et al. [38] proposed the first optimizing compiler for verifiable neural networks using zero-knowledge proofs, named ZEN. ZEN consists of two schemes: ZENacc and ZENinfer. ZENacc proves the accuracy of a committed neural network model, while ZENinfer proves a specific inference result. Together, these schemes ensure both the privacy of user data and the confidentiality of neural network models. To address the computational costs associated with zk-SNARKs, ZEN introduces two optimization algorithms. First, ZEN incorporates a novel neural quantization algorithm that applies two R1CS-friendly optimizations: sign-bit grouping and remainder-based verification. These optimizations enhance the efficiency of converting a floating-point neural network into a fully quantized model expressed in R1CS. As a result, compared to state-of-the-art quantization schemes, ZEN achieves up to 73.9x savings in R1CS constraints for convolution kernels and up to 8.4x reduction for fully connected kernels, all without incurring any additional accuracy loss. Second, ZEN introduces a Single Instruction Multiple Data (SIMD) style optimization, known as stranded encoding. This method optimally encodes multiple low-precision integers (8-bit) with a single finite field element (typically 254-bit). Feng et al. also developed an open-source toolchain that takes a floating-point PyTorch model and converts it into ZEN schemes with all of the aforementioned optimizations. Their evaluation demonstrates that, without any accuracy loss, ZEN provides savings in the number of constraints ranging from 5.43x to 22.19x, with an average reduction of 15.35x, when compared to a vanilla implementation of neural network models in zk-SNARKs.
@@
Liu et al. [10] proposed a scheme for Convolutional Neural Networks (CNNs), called zkCNN, which enables the verification of a model’s inference for a given input data without revealing the model’s parameters. This scheme can also be generalized to prove the model’s accuracy on a public dataset while ensuring that the model’s parameters remain private. zkCNN incorporates a new sumcheck protocol for two-dimensional convolutions, achieving a prover time of $O(n^{2})$ for two input matrices of sizes $n\times n$ and $w\times w$. This is even faster than computing the result directly, making it asymptotically optimal. The proof size is $O(\log n)$. A key component of this protocol is an efficient sumcheck for the Fast Fourier Transform (FFT), which requires only $O(N)$ time to generate the proof for a vector of size $N$, offering a better asymptotic complexity than the conventional $O(N\log N)$ needed for FFT computation. Building upon this, the authors further propose a protocol that achieves a sublinear verifier time of $O(\log^{2}n)$ with a proof size of $O(\log^{2}n)$. Additionally, they design an interactive proof using the Generalized Knowledge Representation protocol for CNN predictions, including verification of the convolutional layers, ReLU activation functions, and max pooling operations. They introduce generalized addition and multiplication gates, which allow operations with fan-in greater than two, enabling inner products to be implemented with a single sumcheck. By extending these gates to accept inputs from any layer, they avoid additional prover time. For CNN convolutional layers, they optimize the sumcheck protocol by reducing the prover time for the inverse FFT (IFFT) by a factor corresponding to the number of input channels. They also design an efficient circuit gadget that combines the computation of the ReLU activation and max pooling functions, requiring only a single bit-decomposition per number. The proposed scheme supports large CNNs like VGG16, which has 15 million parameters and 16 layers, significantly improving the proof generation time to 88.3 seconds—1264 times faster than existing schemes. The proof size is 341 kilobytes, and the verifier time is only 59.3 milliseconds. Furthermore, the scheme can scale to prove the accuracy of the same CNN on 20 images.
Recent advancements in ZKML often underestimate the substantial overhead involved in verifying the necessary commitments to both the model and the data. This challenge becomes particularly pronounced in scenarios such as verifiable testing, where executing the model on larger datasets is required. To address this gap, Ycklama et al. [51] proposed a novel approach for constructing efficient Commit-and-Prove SNARKs (CP-SNARKs) that minimizes the computational burden within the SNARK and extends the underlying proof system in a highly efficient manner. Specifically, they introduce two new CP-SNARK constructions—Apollo and Artemis—that effectively address the emerging challenge of commitment verification in zkML pipelines. Apollo operates on KZG commitments and requires white-box use of the underlying proof system, while Artemis is compatible with any homomorphic polynomial commitment and utilizes only black-box access to the proof system. Apollo simplifies the construction process by eliminating the need for shifting proofs, drastically reducing the number of linking proof instances. This optimization allows Apollo to achieve a 7.3x improvement in prover time compared to Lunar [55]. For instance, when performing an inference proof for MobileNet [56], Lunar requires 20 shifting and 20 linking proofs, whereas Apollo requires just one linking proof. However, Apollo inherits the trusted setup requirement from Lunar. In contrast, Artemis requires only black-box use of the underlying SNARK and supports any homomorphic polynomial commitment, enabling it to work with modern proof systems such as Halo2 with IPA-based commitments, which do not require a trusted setup. The paper also presents the first implementation of Lunar’s CP-SNARK, alongside the implementations of the Apollo and Artemis constructions, all of which are made publicly available as open-source software. The evaluation demonstrates that Apollo and Artemis significantly outperform existing approaches across a range of ML models, improving the state of the art by an order of magnitude. Additionally, the evaluation shows that Artemis, when used without a trusted setup, achieves performance that is effectively the same as Apollo (and Artemis) with a trusted setup.
III-B3 ZKP based Verifiable Inference

### v2

@@ III-B1 ZKP-based Verifiable Training
In 2021, [26] proposed VeriML, whose core idea is to make the training process retrievable to achieve verifiability in machine learning training. VeriML pre-stores the inputs and outputs of several iterations during the training process and commits to them, allowing the prover to retrieve specified iterations upon the verifier’s request and generate proofs of their computational processes. VeriML supports six typical machine learning models, including linear regression, logistic regression, neural networks, SVM, K-Means, and decision trees. Also, VeriML’s computational and communication costs are justified through numerous experiments. Experimental results demonstrate that the communication overhead of VeriML is related to the number of iterations and the number of data samples, and the computational overhead of VeriML is dominated by the batch size of data samples and increases linearly with the batch size.
@@
[45] introduced Hydra, a verifiable training protocol for neural networks built upon the GKR protocol. This system is specifically tailored for neural networks and builds upon the SafetyNets method—the first verifiable inference approach leveraging ZKP, which will be discussed later. In particular, Hydra uses SafetyNets to represent a neural network as an arithmetic circuit. To achieve verifiable training, Hydra introduces a subcircuit protocol and a interactive quantization algorithm. The execution process of Hydra’s subcircuit protocol can be summarized as follows: First, the large and deep circuit representing a neural network is partitioned into multiple smaller subcircuits by depth. The prover then applies the GKR protocol to generate proofs for each sub-circuit. Once a proof for a sub-circuit is generated, it is immediately transmitted to the verifier for verification. Finally, all proofs corresponding to the sub-circuits are aggregated into a single final proof for the entire circuit, which is sent to the verifier for verification. This pipelined approach enables the proof-and-verification process to begin as soon as sub-circuits are uploaded, eliminating the need to process the entire circuit at once. By overlapping proof generation and verification, Hydra significantly enhances verification efficiency. Hydra’s interactive quantization algorithm begins by rounding the weights of the bottom layers of the neural network, freezing these quantized layers, and retraining the network. This process is then repeated layer by layer, progressively moving upward, until all layers have been quantized. This interactive quantization algorithm is executed for each part of the neural network corresponding to each sub-circuit in the pipelined proof-and-verification process.
@@
[27] proposed zkMLaaS, a framework that employs a two-round challenge-response protocol and random sampling of model weights to generate proofs. This approach reduces the time cost of proof generation while ensuring the verifiable integrity of the training process. Before executing the challenge-response protocol, an honest third party (or the ML client) generates a key pair $(\mathsf{PK},\mathsf{VK})$, distributing $\mathsf{PK}$ to the ML service provider $\mathcal{P}$ and $\mathsf{VK}$ to the ML client $\mathcal{V}$. In the first round of the challenge-response protocol, the ML service provider submits commitments for all intermediate weights updated during each training iteration and data sampling epoch to the ML client. In the second round, the ML client randomly selects a subset of these intermediate weights, and the service provider must generate corresponding proofs. The zkMLaaS framework leverages zk-SNARKs and commitment schemes to ensure security. The binding property of the commitment scheme guarantees that the ML service provider cannot produce two different sets of weights corresponding to the same commitment. Furthermore, since the public key is generated by the client or a trusted third party, the service provider is restricted to generating proofs only for the circuits specified by the client. This ensures that the intermediate weights are derived solely from the correct model and data.
@@
zkDL is an efficient ZKML system designed for the verifiable training of deep neural networks [28]. To address the non-arithmetic nature of ReLU activation functions in neural networks, zkDL introduces the zkReLU protocol, which enables efficient proof generation for computations involving ReLU. The zkReLU protocol leverages auxiliary inputs to transform the nonlinear computations of ReLU into an equivalent linear form, enabling efficient handling of forward and backward propagations through ReLU. To ensure the validity of these auxiliary inputs, zkReLU employs the Pedersen commitment scheme and incorporates an anchoring mechanism that links the arithmetic operations within each layer to the auxiliary inputs. This design not only preserves the tensor structure inherent in deep neural networks but also significantly reduces computational redundancy by reusing verified commitments. For circuit design in the training of deep neural networks, zkDL proposes FAC4DNN (Flat Arithmetic Circuit for Deep Neural Networks), which parallelizes the traditionally sequential execution of neural network layers and training iterations. This approach effectively flattens the circuit, reducing its depth by a factor of $O(N)$, where $N$ is the product of the neural network depth and the number of training iterations. zkDL implements a three-step parallel proof process for FAC4DNN: (1) parallel execution of the GKR protocol for each layer, (2) parallel proof of inter-layer arithmetic relationships, and (3) parallel verification of auxiliary inputs. This design ensures that the proof size grows only by $O(\log L)$, where $L$ is the product of the neural network depth. Compared to traditional sequential proof generation, zkReLU’s parallel proof generation offers significant time efficiency. It shows in [28] that the proof generation of zkDL for the entire forward and backward propagation of the neural network training process to be completed within tens of seconds for neural networks with 10 millions parameters.
[29] introduced Zero-Knowledge Proof of Training (zkPoT), a proof-of-training framework aimed at simple models such as logistic regression. Its key idea is to separate arithmetic and non-arithmetic parts of the training computation so that proof size scales with the number of training samples rather than the feature dimension. The design also streams data from auxiliary memory instead of requiring the full computation trace to remain in RAM, making it better aligned with the memory demands of training-oriented workloads.
TABLE VII: Classification of representative ZKML systems by verification objective.
Work Date Model / Scenario Training Testing Inference
@@
MSCProof [50] 2023.10 CNN / image processing $\circ$ $\circ$ $\bullet$
@@
zkGPT [55] 2025.06 LLM $\circ$ $\circ$ $\bullet$
VeriLoRA [31] 2025.08 LLM $\bullet$ $\circ$ $\circ$
While zkPoT [29] demonstrates that proof-of-training can be practical for relatively simple models such as logistic regression, it does not directly address the scale and structure of deep neural network training. To bridge this gap, Abbaszadeh et al. proposed a dedicated proof-of-training framework for DNNs, referred to here as Kaizen [30]. Their design targets small-batch gradient descent and centers on two ideas: an optimized GKR-style proof system for verifying each gradient-descent iteration, and a recursive composition framework for aggregating proofs across many iterations into a single concise proof. At the protocol level, Kaizen tailors a sumcheck-based construction to the algebraic structure of gradient updates, thereby reducing the proving overhead of each step. At the system level, it combines these per-iteration proofs with polynomial commitments to obtain compact proofs and commitments over long training trajectories. This line of work marks an important extension of zkPoT-style thinking from shallow models to modern deep learning workloads.
VeriLoRA extends verifiable training to large language models, with particular emphasis on LoRA-based fine-tuning [31]. Unlike earlier systems that focus mainly on DNNs or CNNs, VeriLoRA offers an end-to-end framework that covers forward propagation, backward propagation, and LoRA weight updates within transformer-based models. For arithmetic-intensive matrix operations, it employs sumcheck protocols over multilinear extensions; for non-arithmetic operations such as SwiGLU, Softmax, element-wise products, and transposition, it adopts specialized lookup-based subprotocols. The system is evaluated on OPT and LLaMA-family models ranging from 3B to 13B parameters. Reported results indicate proving times below 600 seconds per sample, verification times below 4 seconds, and peak GPU memory below 80GB through layer-wise sequential execution. More broadly, VeriLoRA shows that modern parameter-efficient LLM fine-tuning can be brought into the scope of verifiable training without abandoning practical efficiency.
III-B2 ZKP-based Verifiable Testing
@@
In 2020, [32] proposed a ZKML protocol for verifiable inference and accuracy testing of decision trees, named zkDT. zkDT enables the owner of a decision tree model to convince others of the model’s inference results on data samples or its accuracy on public datasets, all without revealing any information about the model itself. The protocol leverages the Aurora protocol [56] as the ZKP backend due to Aurora’s fast proving time, which is a desirable feature for large decision trees. zkDT significantly improves the efficiency of proving time by transforming the computation of decision tree inference into a circuit of size $O(d+h)$, where $d$ is the length of the inference path on the tree, and $h$ is the number of data features. Notably, many testing data samples share common nodes on the decision tree. To exploit this, zkDT optimizes the proof and verification process for decision tree accuracy testing by validating all the nodes of the inference paths across all testing data in a single step, rather than validating the inference path of each sample individually. The implementation of zkDT demonstrates its practicality. For a decision tree with 23 levels and 1029 nodes, and a test dataset consisting of 5000 data samples with 54 features each, zkDT takes 250 seconds to generate a proof of size 287 KB for accuracy testing, and 15.6 seconds for verification.
[52] later revisited decision-tree verification through the lens of lookup arguments and showed how the original zkDT construction can be substantially improved. In general, lookup arguments allow the prover to show that committed values originate from a larger committed table, which is especially attractive for non-arithmetic operations such as range checks, XOR, and AND. Building on this idea, the authors extend vector lookups to matrix lookups, introduce privacy-preserving variants for both the queried substructure and the underlying table, and design more efficient lookup arguments such as cq+, zkcq+, and cq++ [57]. Their application to decision trees is particularly elegant: the tree is encoded as a committed matrix, while the prover commits only to the row corresponding to the reached leaf and then proves that this row belongs to the committed matrix. The prover can subsequently show that the input satisfies the predicates encoded in that row. For batched statements, multiple reached leaves can be handled through a committed matrix of rows. Because the resulting matrix-lookup protocol removes the dependence on tree size from the prover’s asymptotic complexity, it yields a markedly cleaner and more scalable approach to verifiable testing of decision trees.
pvCNN [33] targets verifiable testing of CNNs through a hybrid design that combines homomorphic encryption, collaborative inference, and zk-SNARKs. Its core technical contribution is the use of Quadratic Matrix Programs (QMPs) to encode convolution more efficiently and to support aggregation of proofs across multiple test instances. The framework further splits the model into private and public components, allowing encrypted partial execution while still producing verifiable test results. Relative to earlier QAP-based approaches, pvCNN substantially reduces setup and proving cost for CNN testing.
ZEN [34] can be viewed as one of the first optimizing compilers explicitly designed for verifiable neural networks. It provides two schemes, ZENacc for proving model accuracy and ZENinfer for proving individual inference results, and it improves efficiency mainly through proof-friendly quantization and stranded encoding. These optimizations reduce the R1CS cost of neural operators while preserving model accuracy, making ZEN an important step from hand-crafted circuits toward compiler-assisted ZKML development.
@@
[35] proposed a scheme for Convolutional Neural Networks (CNNs), called zkCNN, which enables the verification of a model’s inference for a given input data without revealing the model’s parameters. This scheme can also be generalized to prove the model’s accuracy on a public dataset while ensuring that the model’s parameters remain private. zkCNN incorporates a new sumcheck protocol for two-dimensional convolutions, achieving a prover time of $O(n^{2})$ for two input matrices of sizes $n\times n$ and $w\times w$. This is even faster than computing the result directly, making it asymptotically optimal. The proof size is $O(\log n)$. A key component of this protocol is an efficient sumcheck for the Fast Fourier Transform (FFT), which requires only $O(N)$ time to generate the proof for a vector of size $N$, offering a better asymptotic complexity than the conventional $O(N\log N)$ needed for FFT computation. Building upon this, the authors further propose a protocol that achieves a sublinear verifier time of $O(\log^{2}n)$ with a proof size of $O(\log^{2}n)$. Additionally, they design an interactive proof using the Generalized Knowledge Representation protocol for CNN predictions, including verification of the convolutional layers, ReLU activation functions, and max pooling operations. They introduce generalized addition and multiplication gates, which allow operations with fan-in greater than two, enabling inner products to be implemented with a single sumcheck. By extending these gates to accept inputs from any layer, they avoid additional prover time. For CNN convolutional layers, they optimize the sumcheck protocol by reducing the prover time for the inverse FFT (IFFT) by a factor corresponding to the number of input channels. They also design an efficient circuit gadget that combines the computation of the ReLU activation and max pooling functions, requiring only a single bit-decomposition per number. The proposed scheme supports large CNNs like VGG16, which has 15 million parameters and 16 layers, significantly improving the proof generation time to 88.3 seconds—1264 times faster than existing schemes. The proof size is 341 kilobytes, and the verifier time is only 59.3 milliseconds. Furthermore, the scheme can scale to prove the accuracy of the same CNN on 20 images.
A separate bottleneck is the cost of checking commitments to models and data inside commit-and-prove pipelines. Apollo and Artemis [36] address this issue by redesigning CP-SNARK linking so that commitment verification places much less burden on the underlying SNARK. Apollo targets KZG-based settings and relies on white-box access to the proof system, whereas Artemis works with generic homomorphic polynomial commitments using only black-box access. This distinction is especially important for ZKML, where commitment checking can dominate the cost of testing or inference proofs on larger models and datasets. In practice, Artemis is particularly attractive because it remains compatible with modern IPA- and Halo2-style settings that avoid trusted setup while still approaching Apollo-level performance.
III-B3 ZKP-based Verifiable Inference